# SHIELD: Secure Hypernetworks for Incremental Expansion Learning Defense

### v2

@@ 4 Enhancing Certified Robustness of SHIELD via Interval MixUp
In this section, we revisit the MixUp technique (Zhang et al. 2018a) and adapt it to align with interval arithmetic and continual learning. We then demonstrate that our interval-based MixUp approach enhances the robustness of our method by encouraging the decision boundary to move as far as possible from intervals to be correctly classified, without sacrificing accuracy.
MixUp
MixUp (Zhang et al. 2018a) improves model generalization by training on convex combinations of input samples and their corresponding labels. Instead of optimizing the model on individual examples, MixUp introduces synthetic points that lie along linear paths between pairs of training samples. This encourages the model to behave smoothly between classes and to avoid overly confident or abrupt predictions in regions where data support is lacking.
MixUp introduces an inductive bias that promotes linear behavior between training examples. It smooths the model’s decision surface, reduces memorization of noisy labels, and improves generalization. By enforcing consistency across interpolated inputs and outputs, it also enhances robustness to ambiguous or uncertain regions in the input space. We present more technical details on MixUp in Appendix B.
Interval MixUp
@@
This formulation yields the maximum radius when $\lambda$ is close to 0 or 1, i.e., when the virtual sample is close to one of the original points, and reduces it to zero when $\lambda=0.5$, i.e., when the sample is equidistant and lies furthest from either original data point. This leads to a smoothed decision boundary and more stable training while maintaining robustness near the data manifold. In Appendix H, we explore the influence of different transformations in Equation (14) (such as linear, quadratic, logarithmic, and cosine ones) on the training process and defense against adversarial attacks.
@@
Furthermore, both the forward and backward passes in Interval MixUp are performed solely on virtual, interval-bounded examples obtained by interpolating pairs of inputs. As a result, the propagated hypercubes (i.e., interval bounds around the inputs) have smaller radii than those centered at the original data points. This has a crucial side effect: it mitigates the wrapping effect (Neumaier 1993), a well-known limitation of interval arithmetic. The wrapping effect arises when high-dimensional intervals are passed through non-linear layers of the neural network, causing them to grow excessively and become increasingly misaligned with the true shape of the reachable set. This over-approximation results in looser, overly conservative bounds that may significantly underestimate the model’s actual robustness.
@@
Figure 2: (left side) Without Interval MixUp: Training leads to sharper boundaries and poor robustness across class transitions. (right side) With Interval MixUp: Interpolated samples ($\times$) with scaled certified regions (boxes) encourage smooth transitions and robust boundaries.
@@
To build intuition for how Interval MixUp operates in practice, we construct a 2D toy experiment, visualized in Figure 2. We generate synthetic training data consisting of a few points from two distinct classes, randomly placed within the unit square $[0,1]^{2}$. Each point is associated with a certified region, shown as a square representing an $\ell_{\infty}$-bounded hypercube.
@@
We then identify several pairs of nearby points belonging to opposite classes. For each such pair, we apply Interval MixUp by linearly interpolating the bounds of their certified regions using a mixing coefficient $\lambda\in[0,1]$. This produces a set of virtual samples along the line connecting each pair. These samples are visualized as $\times$ markers colored by their interpolation ratio, ranging from red (class 0) to blue (class 1), with intermediate hues indicating mixed contributions. Each virtual point is surrounded by a semi-transparent certified box, whose size is adjusted using Equation (14). We train SHIELD model on this setup using only the Interval MixUp samples. The resulting decision surface and certified regions, shown in Figure 2, illustrate how Interval MixUp helps populate the space between classes, guiding the decision boundary away from both the real data.
@@
By interpolating smaller virtual hypercubes, Interval MixUp reduces the growth of these intervals during propagation, thereby maintaining tighter and more informative bounds. This is illustrated in Figure 3, which reports verified accuracy (Zhang et al. 2018b). Verified accuracy measures whether a model is not only correct on a given input but also robust to all perturbations within a specified $\varepsilon$-ball around it, unlike classical accuracy, which only evaluates predictions on clean, unperturbed data. Verified accuracy thus reflects a model’s certified robustness by leveraging the interval bounds computed during training.
@@
Importantly, models trained with Interval MixUp achieve significantly higher verified accuracy compared to those trained with standard IBP, and their verified accuracy closely matches their classical accuracy (see Figure 3). As a consequence, these models outperform standard IBP baselines and are more resilient to adversarial attacks evaluated in this paper. All reported results are based on the best-performing models selected using a single random seed. Moreover, we present examples of Interval MixUp samples in Appendix M.
@@
Figure 3: Comparison of verified accuracy and classical accuracy across four continual learning benchmarks: Permuted MNIST, Rotated MNIST, Split CIFAR-100, and Split miniImageNet. For each task, we report the AA metric after sequentially learning all preceding tasks.
SHIELD Loss Function with Interval MixUp
@@
When applying Interval MixUp in SHIELD, we cannot directly use Equation (9), as the original input data is replaced by virtual samples generated through the interpolation of input pairs. To accommodate this, we introduce a reformulated loss function designed specifically for these synthetic samples.
@@
where $\varepsilon^{\prime}$ is the scaled perturbation radius (see Equation (14)), and $\kappa\in[0,1]$ balances the standard and certified components of the loss. We omit arguments of $\mathcal{L}_{\text{IMixUp}}$ to simplify the notation. The term ^$\hat{\mathcal{L}}_{\text{MixUp}}$ is defined as:
^$\hat{\mathcal{L}}_{\text{MixUp}}$ $\left([\tilde{x}-\varepsilon^{\prime},\tilde{x}+\varepsilon^{\prime}],y_{a},y_{b}\right)=\lambda\cdot\mathcal{L}_{\text{CE}}\left(\sigma\left(\hat{z}_{K,\hat{y}_{a}}\right),y_{a}\right)$ (17)
@@
where $\hat{z}_{K,\hat{y}_{a}}$ and $\hat{z}_{K,\hat{y}_{b}}$ denote the components of the target network’s interval output associated with classes $\hat{y}_{a}$ and $\hat{y}_{b}$, respectively, after $K$ transformations, as specified in Equation (8). This formulation enables us to retain certified robustness guarantees even for interpolated virtual samples, allowing us to benefit from MixUp’s generalization while maintaining interval-based certification.
Finally, to adapt Interval MixUp to the continual learning setting, we incorporate it into the overall objective by combining it with the hypernetwork regularization term from Equation (10). The complete loss function becomes:
@@
where $\beta$ is a hyperparameter controlling the strength of the hypernetwork regularization, and $\mathcal{L}_{\text{out}}$ is defined as in Equation (10). This final formulation ensures that our model can learn robust, certified representations in the context of continual learning. The training algorithm of our method is provided in Appendix K.
@@ 5 Experiments
This section outlines the experimental setup and results for SHIELD. We describe the datasets used, along with the architectures of both the target network and the hypernetwork. All experiments are conducted in a Task-Incremental Learning setting, where the task identity is known to the model during both training and evaluation. Moreover, our method naturally extends to the Class-Incremental Learning (CIL) scenario, retaining a significant level of adversarial robustness. Appendix J shows that SHIELD consistently outperforms a strong baseline under various attacks in the CIL scenario.
Benchmarks
To assess the effectiveness of our approach, we perform experiments on four standard benchmarks: Permuted MNIST, Rotated MNIST, Split CIFAR-100, and Split miniImageNet. The structure of tasks and class divisions follows the protocol established in (Ru et al. 2024). Full dataset and task details are provided in Appendix D.
Architectures
For Permuted MNIST, Rotated MNIST, and Split miniImageNet, we adopt the same target network architectures as in (Ru et al. 2024), with key modifications: we introduce interval-aware variants of normalization and pooling layers, include biases, and discard the use of separate convolutional layers per task. For Split CIFAR-100, we use a streamlined interval-based version of AlexNet. Further architectural details are provided in Appendix G.
Metrics
@@
Results
Table 1 reports the AA and BWT obtained by SHIELD and several baselines after sequential learning over all tasks in the Permuted MNIST, Rotated MNIST, Split CIFAR-100, and Split miniImageNet.
Table 1: AA performance after all tasks on the Permuted/Rotated MNIST, Split CIFAR-100, and Split miniImageNet. Results for SHIELD and SHIELD with Interval MixUp (SHIELDIM) are averaged over 2 seeds for AutoAttack and 5 seeds for the other evaluations, except for Split miniImageNet (3 seeds, 2 for AutoAttack). Baseline results are from (Ru et al. 2024). Extended results with standard deviations are in Appendix I
@@ Results
As shown in Table 1, on Split miniImageNet, SHIELD significantly outperforms all baselines across all metrics. While DGP and GPM reach original sample AA of 44.8% and 36.8%, respectively, SHIELD achieves 59.52%, and SHIELDIM reaches 62.67%. It achieves the best AA on every task and the overall average. Forgetting (measured by BWT) is also well-controlled, with SHIELD and SHIELDIM reaching -0.16 and -0.18, respectively, while achieving far higher AA.
@@ 6 Conclusions
SHIELD is the first certifiably robust continual learning method, combining hypernetworks with interval arithmetic to ensure robustness across tasks. Its core, Interval MixUp, assigns smaller perturbation radii to virtual points farther from real data, tightening bounds and pushing decision boundaries away from certified regions, enhancing both robustness and accuracy, even under large perturbations and on challenging datasets.
Furthermore, SHIELD represents the first demonstrated approach to achieving substantial certified robustness in the CIL setting. This highlights the method’s ability to unify continual adaptation with formal guarantees, marking an important step toward robust lifelong learning.
Limitations While SHIELD achieves strong certified robustness and adaptability across tasks, the use of IBP, despite its computational efficiency, may lead to overly conservative (i.e., wide) interval bounds. This can affect both the certified robustness and the efficiency of the continual learner, especially in high-dimensional settings.
Appendix A Extended related works
@@
Adversarial attacks on deep neural networks are typically categorized as white-box or black-box, depending on the attacker’s access to the target model. In white-box settings, the adversary fully knows the model, including its architecture, parameters, and gradients. Prominent white-box attacks include gradient-based methods such as FGSM (Goodfellow, Shlens, and Szegedy 2014), BIM (Huang et al. 2011), MIM (Dong et al. 2018), PGD (Madry et al. 2017), and AutoAttack (De Lange et al. 2021), as well as optimization-based approaches like the OnePixel (Su, Vargas, and Sakurai 2019).
@@
In black-box settings, attackers have limited access to the model and must rely on indirect signals. Some approaches use confidence scores to estimate gradients, such as zoo (Chen et al. 2017), while others operate with only the predicted label, as seen in decision-based methods like the Boundary Attack (Brendel, Rauber, and Bethge 2017). Another strategy involves crafting adversarial examples on a surrogate model and transferring them to the target model, exploiting the cross-model transferability of adversarial perturbations (Dong et al. 2018, 2019).
@@
Simultaneously with adversarial attack techniques, there are many defence algorithms. Early adversarial defenses relied on heuristics such as input transformations (Xie et al. 2017), model ensembles (Bagnall, Bunescu, and Stewart 2017), and denoisers (Jin et al. 2019), but many were later found to rely on obfuscated gradients (Athalye, Carlini, and Wagner 2018). More robust approaches like adversarial training (AT) (Goodfellow, Shlens, and Szegedy 2014; Jia et al. 2022) and defensive distillation (Goldblum et al. 2020; Zhu et al. 2021) have become dominant. Recent work has enhanced AT with learnable strategies (Jia et al. 2022), feature perturbation (Mustafa et al. 2020), and optimization techniques (Pang et al. 2020). However, the above models are typically static and struggle against evolving attack sequences. Test-Time Adaptation Defense (Shi, Holtz, and Mishne 2021; Yang et al. 2022) addresses adaptation to new attacks but neglects previous ones.
@@
Most existing methods aim to improve adversarial robustness, but they do so without providing formal guarantees. Interval Bound Propagation (IBP) (Gowal et al. 2018; Mirman, Gehr, and Vechev 2018) offers a principled alternative by using interval arithmetic to train models that are provably robust against adversarial attacks. These methods enforce that the model classifies all points within a specified perturbation region (typically a ball in $\ell_{\infty}$) correctly, by optimizing a worst-case cross-entropy loss.
Continual Learning
@@
Continual learning methods are often grouped into architectural, rehearsal, and regularization-based strategies (Hsu et al. 2018). Architectural approaches like Progressive Neural Networks (Rusu et al. 2016) and CopyWeights with Re-init (Lomonaco and Maltoni 2017) prevent forgetting by expanding or modifying network structures. Further developments reuse a fixed architecture with iterative pruning, such as PackNet (Mallya and Lazebnik 2018) or Supermasks (Wortsman et al. 2020). Rehearsal-based methods, such as the buffer strategy in (Hayes, Cahill, and Kanan 2019), retain selected past examples, while pseudo-rehearsal approaches train generative models to replay prior data (Mazur et al. 2022). Regularization-based techniques like EWC (Kirkpatrick et al. 2017), SI (Zenke, Poole, and Ganguli 2017), LwF (Li and Hoiem 2017), and MAS (Aljundi et al. 2018) constrain parameter updates based on past task importance, often using the Fisher Information Matrix or similar metrics.
@@
Hypernetworks, introduced in (Ha, Dai, and Le 2016), are defined as neural models that generate weights for a separate target network that solves a specific task. The authors aim to reduce the number of trainable parameters by designing a hypernetwork that often has fewer parameters than the target network. In continual learning, hypernetworks can directly generate individual weights for subsequent continual learning tasks, like in HNET (von Oswald et al. 2019). Then, HNET was extended with a probabilistic approach called posterior meta-replay (Henning et al. 2021). In HyperMask (Książek and Spurek 2023), the authors used the lottery ticket hypothesis to produce task-specific masks for the trainable or fixed target model. In (Krukowski et al. 2024), the authors employ interval arithmetic on weights to regularize the target model. These methods are positioned between those based on architecture and those focused on regularization.
@@
The above concepts have many different modifications and upgrades (Wang et al. 2024). But the adversarial robustness of continual learning problems is unexplored in the literature. AIR (Zhou and Hua 2024) introduced the first framework for continual adversarial defense, addressing robustness against evolving attack sequences. The method leverages unsupervised data augmentation to enhance general robustness in a task-agnostic manner. While innovative, AIR is limited in scalability, having been evaluated on only two to three defense tasks, raising concerns about its effectiveness in longer, continual settings.
@@
Alternatively, the study on Double Gradient Projection (DGP) (Ru et al. 2024) highlights that this technique is notably more resilient to adversarial attacks compared to traditional continual learning approaches such as GEM (Lopez-Paz and Ranzato 2017) and GPM (Saha, Garg, and Roy 2021). DGP enhances adversarial robustness by projecting gradients orthogonally concerning a critical subspace before updating weights, thus maintaining the prior gradient smoothness from earlier samples. Furthermore, (Khan et al. 2022) examines the resistance of current models against adversarial assaults. These methodologies enhance the robustness of continual learning, yet they have several significant limitations. Firstly, such approaches are restricted to several subsequent tasks and do not give strict guarantees.
Appendix B MixUp
Definition B.1 (MixUp).
Given two training examples $(x_{i},y_{i})$ and $(x_{j},y_{j})$ drawn independently from the data distribution, MixUp constructs a new virtual training example $(\tilde{x},\tilde{y})$ as:

### v3

@@ 4 Enhancing Certified Robustness of SHIELD via Interval MixUp
In this section, we revisit the MixUp technique [zhang2018mixupempiricalriskminimization] and adapt it to align with interval arithmetic and continual learning. We then demonstrate that our interval-based MixUp approach enhances the robustness of our method by encouraging the decision boundary to move as far as possible from intervals to be correctly classified, without sacrificing accuracy.
MixUp.
MixUp [zhang2018mixupempiricalriskminimization] improves model generalization by training on convex combinations of input samples and their corresponding labels. Instead of optimizing the model on individual examples, MixUp introduces synthetic points that lie along linear paths between pairs of training samples. This encourages the model to behave smoothly between classes and to avoid overly confident or abrupt predictions in regions where data support is lacking.
MixUp introduces an inductive bias that promotes linear behavior between training examples. It smooths the model’s decision surface, reduces memorization of noisy labels, and improves generalization. Enforcing consistency across interpolated inputs and outputs also enhances robustness to ambiguous or uncertain regions in the input space. Additional technical details on MixUp are provided in SM 8.
Interval MixUp.
@@
This formulation yields the maximum radius when $\lambda$ is close to 0 or 1, i.e., when the virtual sample is close to one of the original points, and reduces it to zero when $\lambda=0.5$, i.e., when the sample is equidistant and lies furthest from either original data point. This leads to a smoothed decision boundary and more stable training while maintaining robustness near the data manifold. In SM 14, we explore the influence of different transformations in Equation (15) (such as linear, quadratic, logarithmic, and cosine ones) on the training process and defense against adversarial attacks.
@@
Furthermore, both the forward and backward passes in Interval MixUp are performed solely on virtual, interval-bounded examples obtained by interpolating pairs of inputs. As a result, the propagated hypercubes (i.e., interval bounds around the inputs) have smaller radii than those centered at the original data points. This has a crucial side effect: it mitigates the wrapping effect [neumaier1993wrapping], a well-known limitation of interval arithmetic. The wrapping effect arises when high-dimensional intervals are passed through non-linear layers of the neural network, causing them to grow excessively and become increasingly misaligned with the true shape of the reachable set. This over-approximation results in looser, overly conservative bounds that may significantly underestimate the model’s actual robustness.
@@
Rysunek 2: (left side) Without Interval MixUp: Training leads to sharper boundaries and poor robustness across class transitions. (right side) With Interval MixUp: Interpolated samples ($\times$) with scaled certified regions (boxes) encourage smooth transitions and robust boundaries.
@@
To build intuition for how Interval MixUp operates in practice, we construct a 2D toy experiment, visualized in Fig. 2. We generate synthetic training data consisting of a few points from two distinct classes, randomly placed within the unit square $0,1]^{2}$. Each point is associated with a certified region, shown as a square representing an $\ell_{\infty}$-bounded hypercube.
@@
We then identify several pairs of nearby points belonging to opposite classes. For each such pair, we apply Interval MixUp by linearly interpolating the bounds of their certified regions using a mixing coefficient $\lambda\in[0,1]$. This produces a set of virtual samples along the line connecting each pair. These samples are visualized as $\times$ markers colored by their interpolation ratio, ranging from red (class 0) to blue (class 1), with intermediate hues indicating mixed contributions. Each virtual point is surrounded by a semi-transparent certified box, whose size is adjusted using Eq. (15). We train SHIELD model on this setup using only the Interval MixUp samples. The resulting decision surface and certified regions, shown in Fig. 2, illustrate how Interval MixUp helps populate the space between classes, guiding the decision boundary away from both the real data.
@@
By interpolating smaller virtual hypercubes, Interval MixUp reduces the growth of these intervals during propagation, thereby maintaining tighter and more informative bounds. This is illustrated in Fig. 3, which reports verified accuracy [zhang2018efficientneuralnetworkrobustness]. Verified accuracy measures whether a model is not only correct on a given input but also robust to all perturbations within a specified $\varepsilon$-ball around it, unlike classical accuracy, which only evaluates predictions on clean, unperturbed data. Verified accuracy thus reflects a model’s certified robustness by leveraging the interval bounds computed during training.
@@
Importantly, models trained with Interval MixUp achieve significantly higher verified accuracy compared to those trained with standard IBP, and their verified accuracy closely matches their classical accuracy (see Figure 3). As a consequence, these models outperform standard IBP baselines and are more resilient to adversarial attacks evaluated in this paper. All reported results are based on the best-performing models selected using a single random seed. Moreover, we present examples of Interval MixUp samples in SM 20.
@@
Rysunek 3: Comparison of verified accuracy and classical accuracy across four continual learning benchmarks: Permuted MNIST, Rotated MNIST, Split CIFAR-100, and Split miniImageNet. For each task, we report the AA metric after sequentially learning all preceding tasks.
SHIELD Loss Function with Interval MixUp.
@@
When applying Interval MixUp in SHIELD, we cannot directly use Eq. (9), as the original input data is replaced by virtual samples generated through the interpolation of input pairs. To accommodate this, we introduce a reformulated loss function designed specifically for these synthetic samples.
@@
where $\varepsilon^{\prime}$ is the scaled perturbation radius (see Eq. (15)), and $\kappa\in[0,1]$ balances the standard and certified components of the loss. We omit arguments of $\mathcal{L}_{\text{IMixUp}}$ to simplify the notation. The term $\hat{\mathcal{L}}_{\text{MixUp}}$ is defined as:
$\hat{\mathcal{L}}_{\text{MixUp}}$ $\left([\tilde{x}-\varepsilon^{\prime},\tilde{x}+\varepsilon^{\prime}],y_{a},y_{b}\right)=\lambda\cdot\mathcal{L}_{\text{CE}}\left(\sigma\left(\hat{z}_{K,\hat{y}_{a}}\right),y_{a}\right)$ (18)
@@
where $\hat{z}_{K,\hat{y}_{a}}$ and $\hat{z}_{K,\hat{y}_{b}}$ denote the components of the target network’s interval output associated with classes $\hat{y}_{a}$ and $\hat{y}_{b}$, respectively, after $K$ transformations, as specified in Eq. (8). This formulation enables us to retain certified robustness guarantees even for interpolated virtual samples, allowing us to benefit from MixUp’s generalization while maintaining interval-based certification.
Finally, to adapt Interval MixUp to the continual learning setting, we incorporate it into the overall objective by combining it with the hypernetwork regularization term from Eq. (10). The complete loss function becomes:
@@
where $\beta$ is a hyperparameter controlling the strength of the hypernetwork regularization, and $\mathcal{L}_{\text{out}}$ is defined as in Eq. (10). This final formulation ensures that our model can learn robust, certified representations in the context of continual learning. The training algorithm of our method is provided in SM 18.
@@ 5 Experiments
This section presents the experimental setup and results for SHIELD. We first describe the datasets and the architectures of both the target network and the hypernetwork. All experiments are conducted in the Task-Incremental Learning setting, where the task identity is provided to the model during both training and evaluation. In addition, our method naturally extends to the Class-Incremental Learning (CIL) scenario while retaining a significant degree of adversarial robustness. SM 16 shows that SHIELD consistently outperforms a strong baseline under multiple attack settings in the CIL scenario.
Since AIR [zhou2024defense] is not attack-agnostic and therefore operates under a different set of assumptions than SHIELD, we include a separate comparison in SM 17. Notably, SHIELD outperforms AIR despite being task-agnostic, whereas AIR is not.
Benchmarks.
To assess the effectiveness of our approach, we perform experiments on four standard benchmarks: Permuted MNIST, Rotated MNIST, Split CIFAR-100, and Split miniImageNet. The structure of tasks and class divisions follows the protocol established in [ru2024maintaining]. Full dataset and task details are provided in SM 10.
Architectures.
For Permuted MNIST, Rotated MNIST, and Split miniImageNet, we adopt the same target network architectures as in [ru2024maintaining], with key modifications: we introduce interval-aware variants of normalization and pooling layers, include biases, and discard the use of separate convolutional layers per task. For Split CIFAR-100, we use a streamlined interval-based version of AlexNet. Further architectural details are provided in SM 13.
Metrics.
@@
Results.
Tab. 1 reports the AA and BWT obtained by SHIELD and several baselines after sequential learning over all tasks in the Permuted MNIST, Rotated MNIST, Split CIFAR-100, and Split miniImageNet.
Tabela 1: AA performance after all tasks on the Permuted/Rotated MNIST, Split CIFAR-100, and Split miniImageNet. Results for SHIELD and SHIELD with Interval MixUp (SHIELDIM) are averaged over 2 seeds for AutoAttack and 5 seeds for the other evaluations, except for Split miniImageNet (3 seeds, 2 for AutoAttack). Baseline results are from [ru2024maintaining]. Extended results with standard deviations are in SM 15
@@ Results.
As shown in Tab. 1, on Split miniImageNet, SHIELD significantly outperforms all baselines across all metrics. While DGP and GPM reach original sample AA of 44.8% and 36.8%, respectively, SHIELD achieves 59.52%, and SHIELDIM reaches 62.67%. It achieves the best AA on every task and the overall average. Forgetting (measured by BWT) is also well-controlled, with SHIELD and SHIELDIM reaching -0.16 and -0.18, respectively, while achieving far higher AA.
@@ 6 Conclusions
SHIELD is the first certifiably robust continual learning method, combining hypernetworks with interval arithmetic to ensure robustness across tasks. Its core component, Interval MixUp, assigns smaller perturbation radii to virtual points farther from real data, tightening bounds and pushing decision boundaries away from certified regions to enhance robustness and accuracy even under large perturbations.
Moreover, SHIELD is the first approach to demonstrate substantial certified robustness in the CIL setting, showing that continual adaptation and formal guarantees can be unified, an important step toward robust lifelong learning.
Limitations.
While SHIELD achieves strong certified robustness and adaptability, the use of IBP, though efficient, may yield overly conservative (i.e., wide) output bounds. This can affect both robustness and computational efficiency, particularly in high-dimensional settings.
[…]
Supplementary Material
7 Extended related works
@@
Adversarial attacks on deep neural networks are typically categorized as white-box or black-box, depending on the attacker’s access to the target model. In white-box settings, the adversary fully knows the model, including its architecture, parameters, and gradients. Prominent white-box attacks include gradient-based methods such as FGSM [goodfellow2014explaining], BIM [huang2011adversarial], MIM [dong2018boosting], PGD [madry2017towards], and AutoAttack [de2021continual], as well as optimization-based approaches like the OnePixel [su2019one].
@@
In black-box settings, attackers have limited access to the model and must rely on indirect signals. Some approaches use confidence scores to estimate gradients, such as zoo [chen2017zoo], while others operate with only the predicted label, as seen in decision-based methods like the Boundary Attack [brendel2017decision]. Another strategy involves crafting adversarial examples on a surrogate model and transferring them to the target model, exploiting the cross-model transferability of adversarial perturbations [dong2018boosting, dong2019evading].
@@
Simultaneously with adversarial attack techniques, there are many defence algorithms. Early adversarial defenses relied on heuristics such as input transformations [xie2017mitigating], model ensembles [bagnall2017training], and denoisers [jin2019ape], but many were later found to rely on obfuscated gradients [athalye2018obfuscated]. More robust approaches like adversarial training (AT) [goodfellow2014explaining, jia2022adversarial] and defensive distillation [goldblum2020adversarially, zhu2021reliable] have become dominant. Recent work has enhanced AT with learnable strategies [jia2022adversarial], feature perturbation [mustafa2020deeply], and optimization techniques [pang2020bag]. However, the above models are typically static and struggle against evolving attack sequences. Test-Time Adaptation Defense [shi2021online, yang2022adaptive] addresses adaptation to new attacks but neglects previous ones.
@@
Most existing methods aim to improve adversarial robustness, but they do so without providing formal guarantees. Interval Bound Propagation (IBP) [gowal2018effectiveness, mirman2018differentiable] offers a principled alternative by using interval arithmetic to train models that are provably robust against adversarial attacks. These methods enforce that the model classifies all points within a specified perturbation region (typically a ball in $\ell_{\infty}$) correctly, by optimizing a worst-case cross-entropy loss.
Continual Learning.
@@
Continual learning methods are often grouped into architectural, rehearsal, and regularization-based strategies [hsu2018re]. Architectural approaches like Progressive Neural Networks [rusu2016progressive] and CopyWeights with Re-init [lomonaco2017core50] prevent forgetting by expanding or modifying network structures. Further developments reuse a fixed architecture with iterative pruning, such as PackNet [mallya2018packnet] or Supermasks [wortsman2020supermasks]. Rehearsal-based methods, such as the buffer strategy in [hayes2019memory], retain selected past examples, while pseudo-rehearsal approaches train generative models to replay prior data [mazur2022target]. Regularization-based techniques like EWC [kirkpatrick2017overcoming], SI [zenke2017continual], LwF [li2017learning], and MAS [aljundi2018memory] constrain parameter updates based on past task importance, often using the Fisher Information Matrix or similar metrics.
@@
Hypernetworks, introduced in [ha2016hypernetworks], are defined as neural models that generate weights for a separate target network that solves a specific task. The authors aim to reduce the number of trainable parameters by designing a hypernetwork that often has fewer parameters than the target network. In continual learning, hypernetworks can directly generate individual weights for subsequent continual learning tasks, like in HNET [von2019continual]. Then, HNET was extended with a probabilistic approach called posterior meta-replay [henning2021posterior]. In HyperMask [ksikazek2023hypermask], the authors used the lottery ticket hypothesis to produce task-specific masks for the trainable or fixed target model. In [krukowski2024hyperinterval], the authors employ interval arithmetic on weights to regularize the target model. These methods are positioned between those based on architecture and those focused on regularization.
@@
The above concepts have many different modifications and upgrades [wang2024comprehensive]. But the adversarial robustness of continual learning problems is unexplored in the literature. AIR [zhou2024defense] introduced the first framework for continual adversarial defense, addressing robustness against evolving attack sequences. The method leverages unsupervised data augmentation to enhance general robustness in a task-agnostic manner. While innovative, AIR is limited in scalability, having been evaluated on only two to three defense tasks, raising concerns about its effectiveness in longer, continual settings.
@@
Alternatively, the study on Double Gradient Projection (DGP) [ru2024maintaining] highlights that this technique is notably more resilient to adversarial attacks compared to traditional continual learning approaches such as GEM [Paz2017gradient] and GPM [saha2021gradient]. DGP enhances adversarial robustness by projecting gradients orthogonally concerning a critical subspace before updating weights, thus maintaining the prior gradient smoothness from earlier samples. Furthermore, [khan2022susceptibility] examines the resistance of current models against adversarial assaults. These methodologies enhance the robustness of continual learning, yet they have several significant limitations. Firstly, such approaches are restricted to several subsequent tasks and do not give strict guarantees.
8 MixUp
Definition 8.1 (MixUp).
[…]